BLTWG Meeting 03/02/2015
- Time: 11:30 EST
- Connecting: detailed instructions are here
- H.323 ip address: 184.108.40.206
- web browser: https://bluejeans.com/536076821
- phone: 408-740-7256 or 888-240-2560
- Conference id: 536076821
- Present: Alex B, Nathan S., Kei M., Tim W., Alexandre D., Eugene C., Curtis M., James M., Brendan P., Franz K., Richard J., Alex S.
- Update on polarimetry - Kei M.
- Update on the fixed tagger hodoscope - Nathan S.
- Update on the pair spectrometer - Alex S.
- Update on the tagger microscope - Alex B.
- Update on diamond radiators - Brendan P.
- inventory of GlueX diamonds
- report from CHESS run: Feb. 18-24
- diamond contract status
- any other business
Fast feedback will not be ready for the run. No expected completion date. Magnet response hasn't been checked.
Work is continuing with the polarimeter.
Nathan has found signal cable issues in highest electron energy counters of TAGH. Looking into it.
Alex B. turned on TAGM to check power. Some control boards have wrong read back values. Looks like readback resistors not making proper contact.
Brendan is looking into epoxies for mounting diamonds and is continuing the analysis of CHESS data. Alex B. and Nathan S. gave Tim the diamond mount with 3 diamonds. Hoop will be mounted and surveyed tomorrow (Tuesday). Sinmat is unable to thin diamonds as originally proposed. One possibility is to use a laser to cleave a diamond into 2 with help from Applied Diamond, or to purchase diamonds from a company in Singapore.
Alex S. is busy with alignment of coarse counters.
Alex B. will rearrange the preamps to put noisy ones on the outer wedges.